OpenStudy (anonymous):

Nina is 10 years younger than Deepak. Deepak is 3 times as old as Nina. Which system of equations can be used to find d, Deepak's age, and n, Nina's age? d = n + 10 d = 3n d = n − 10 d = 3n d = n + 10 n = 3d d = n − 10 n = 3d

OpenStudy (chrisdbest):

I would go with the first one

OpenStudy (anonymous):

I was thinking the same thing... I just wasn't really sure... It's confusing me >.<

OpenStudy (chrisdbest):

We know Nina is 10 years younger than my niqqa Deepak, so D = 10 - N, or D = N + 10.

OpenStudy (chrisdbest):

And Deepak is 3 times older than Nina, so D = 3N

OpenStudy (anonymous):

xD Done!

OpenStudy (anonymous):

D= 10